Solve for the determinant in the equation below.

4 −3 2

1 2 −2

2 -1 −4


1
Expert's answer
2022-03-15T19:25:45-0400

Solution


Given that


"det(A)=\begin{vmatrix}
 4 & -3 & 2\\
 1 & 2 & -2\\
 2 & -1 & -4
\end{vmatrix}"


"det(A)=4\begin{vmatrix}
 2 & -2 \\
 -1 & -4
\end{vmatrix}-(-3)\begin{vmatrix}
 1 & -2 \\
 2 & -4
\end{vmatrix}+2\begin{vmatrix}
 1 & 2 \\
 2 & -1
\end{vmatrix}"


"det(A)=4(-8-2)+3(-4+4)+2(-1-4)"



"det(A)=" "4(-10)+3(0)+2(-5)"


"det(A)=-40+0-10"


"det(A)=-50"